OpenStudy (rawr7548):

simplify What is 4/5 divded by 1/8 divided by 1/2 show all work

OpenStudy (snowfire):

When you divide, you're actually just multiplying by the reciprocal, so dividing by 1/8 is just multiplying by 8/1=8. Think you can handle the problem now?

OpenStudy (rawr7548):

No can you do the rest please

OpenStudy (snowfire):

I'll do the first part, since you need to actually do it to solidify your understanding. (4/5)/(1/8)=(4/5)*8=32/5

OpenStudy (rawr7548):

Okay is that all of the work? or there is more to the problem

OpenStudy (snowfire):

That's the first operation, now tell me what you get when you divide that by 1/2

OpenStudy (rawr7548):

32/5 divdied by 1/2

OpenStudy (snowfire):

And what does that give you?